OpenStudy (anonymous):

Solve for all values. sinx + sqrt3 = -sinx

OpenStudy (anonymous):

\[\sin x +\sqrt{3} = -\sin x\]

OpenStudy (anonymous):

Use algebra to solve for what you want to know. Solving for something with algebra means that you isolate all occurrences of it on one side of the equals sign. 1) Your unknown is sin(x), so get all the sin(x)'s alone on one side of the equals sign. 2) Now undo the sin() function by applying the inverse function on both sides of the equation. -Kyle

OpenStudy (mathstudent55):